OpenStudy (anonymous):

Write a point-slope equation for the line that has slope 13 and passes through the point (15, 12)

OpenStudy (anonymous):

Ok so remember that point-slope formula is: y-y1=m(x-x1) ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok so do u remember what the M stands 4 in the equation?

OpenStudy (anonymous):

slope

OpenStudy (anonymous):

yeah and what's the slope in this problem? :)

OpenStudy (anonymous):

13

OpenStudy (anonymous):

perfect! Ok now do u remember what the x1 & y1 are?

OpenStudy (anonymous):

15 and 12

OpenStudy (anonymous):

yep nice ok, so now try to put the numbers into their places and show me what u get :)

OpenStudy (anonymous):

y-12=13(x-15)

OpenStudy (anonymous):

perfect! Ok is that as far as they wanted u 2 go or they wanted it in slope-intercept form? (y=mx+b) ?

OpenStudy (anonymous):

point slope equation

OpenStudy (anonymous):

Oh ok so I'm pretty sure that's it then :) is that one of ur answers?

OpenStudy (anonymous):

had to type it in but yes it was right thanks

OpenStudy (anonymous):

ok perfect! :) getting 2 understand how 2 do them a bit better?

OpenStudy (anonymous):

yea i just have to have some check my aswerns to make sure there right

OpenStudy (anonymous):

cool ok yeah I'll help u with all that u need

OpenStudy (anonymous):

question 2 of 10 lol il post a new question

OpenStudy (anonymous):

k np :)